# Measuring Data Science Automation: A Survey of Evaluation Tools for AI Assistants and Agents

### v1

@@ Abstract
Data science aims to extract insights from data to support decision-making processes. Recently, Large Language Models (LLMs) are increasingly used as assistants for data science, by suggesting ideas, techniques and small code snippets, or for the interpretation of results and reporting. Proper automation of some data-science activities is now promised by the rise of LLM agents, i.e., AI systems powered by an LLM equipped with additional affordances—such as code execution and knowledge bases—that can perform self-directed actions and interact with digital environments. In this paper, we survey the evaluation of LLM assistants and agents for data science. We find (1) a dominant focus on a small subset of goal-oriented activities, largely ignoring data management and exploratory activities; (2) a concentration on pure assistance or fully autonomous agents, without considering intermediate levels of human-AI collaboration; and (3) an emphasis on human substitution, therefore neglecting the possibility of higher levels of automation thanks to task transformation.
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020) and their multimodal extensions first caught public prominence by powering capable chatbots that are now widely used as assistants to humans in several tasks, such as summarising documents (Liu et al., 2023c), performing translations (Zhu et al., 2023), and creating code snippets (Guo et al., 2023). These LLM assistants take instructions from a human in the form of a prompt and return an answer, with the human retaining control over planning and decision-making by determining the sequence of actions to follow and how much to rely on the assistant’s output. Attention is now increasingly dedicated to “LLM agents(Wang et al., 2024) that can autonomously and iteratively decide a sequence of actions to take and repeatedly interact with an external (digital) environment, being equipped with affordances and tools such as code execution (Huang et al., 2024b), internet access (Zhou et al., 2024), knowledge bases (Chen et al., 2024), and operating system control (Liu et al., 2023a).
@@
In this paper we focus on how to evaluate assistants and agents for data science applications. Data science is the process of handling and analysing data to extract insights that support decision-making in science, business, or other contexts. Data science may deal with different modalities of data (tabular, images, audio, etc.) but it always involves writing and interpreting information represented as text, whether in the form of code or natural language, and processing image, such as to present information or models. This dominance of textual modalities, combined with the vast amount of relevant online material that LLMs leverage during both training and inference, makes data science well-suited for automation by LLMs. Indeed, since the early days of LLMs (Chandel et al., 2022), LLM assistants and agents have been used in data science applications. In this work, we survey evaluation tools measuring performance on tasks across the data science pipeline, offering, to our knowledge, the first comprehensive review of LLM evaluation for core data science tasks [^1] While many studies assess LLMs on tasks related to data science, such as coding (Jimenez et al., 2024) and planning (Valmeekam et al., 2023), our focus is on those that explicitly target data science.. We do not overview the state of the art in developing LLM assistants or agents for data science, referring interested readers to other works (Sun et al., 2024).
@@
Data science is highly multidisciplinary and involves a breadth of activities, combining fields such as statistics, machine learning, and data engineering with tasks such as understanding business needs, writing reports, and preliminary research. Therefore, to compare the data science automation evaluation tools in our survey, we adopt the widely used data science task taxonomy of Martínez-Plumed et al. (2019) and classify each evaluation tool by the activities it requires subjects to perform and explicitly evaluates. Moreover, in our survey we specifically consider the level of autonomy each evaluation targets—whether the subjects are agents, assistants, or intermediate forms, such as LLM agents operating under close human supervision that correct their actions as needed. Further, we also analyse the way in which the tasks are framed and evaluated, to understand if they measure the ability of AI systems to simply substitute humans or if instead they consider that the AI system can transform the task in deeper ways, such as by bringing functional improvements (for instance, LLMs may not need to produce visualisations to perform data exploration; see Sec. 2.1). By focusing on activities and autonomy, we make the following findings:
@@
  • Most evaluation works individually target a (small) subset of data science activities (Tables 2 and 3); a few works cover multiple activities (Secs. 4.2 and 4.3), but none cover all of them. Taken together, the surveyed works cover the landscape of data-science activities in a biased fashion, chiefly over-representing “goal-oriented” activities, such as data preprocessing, producing plots in specified formats, or building predictive models for predetermined targets. Only a few studies (Cheng et al., 2023; Sahu et al., 2025; Majumder et al., 2024) give prominence to open-ended, exploratory aspects of data science (such as interpreting client needs within a business context, creatively exploring datasets and proposing potential uses) or data management(Yu et al., 2018; 2019b; 2019a; Lei et al., 2024).
@@
  • Most studies focus either on assistants following human-defined actions or fully autonomous agents, overlooking more realistic scenarios of intermediate LLM–human collaboration. Exceptions include Li et al. (2024; 2025b; 2025a) (agents with simulated human users) and Yu et al. (2019a) (assistants aiding users in clarifying data tasks). Due to this near-binary focus, we organise the surveyed works according to the focus on assistants (Sec. 3) and agents (Sec. 4), highlighting when they assess intermediate autonomy.
@@
  • Evaluations often implicitly assume that AI will substitute humans without functionally changing the tasks, either in assuming the steps by which a task is solved are the same a human would follow (Yu et al., 2019a; Zhang et al., 2024b; Chen et al., 2024) or by scoring the task referring to human-produced output, despite there not being a single ground truth (Song et al., 2025; Huang et al., 2024b; Hu et al., 2024; Jing et al., 2024; Chen et al., 2024). Other works instead reward agents that transform the task to solve it by scoring the final output (Pietruszka et al., 2024; Cheng et al., 2023; Li et al., 2025b; Chan et al., 2024; Lei et al., 2024; Majumder et al., 2024; Sahu et al., 2025).
@@ 2.3 LLM evaluation
The area of AI evaluation (Burden et al., 2025) mostly relies on tasks encapsulated in input-output benchmarks with a reference output for each example. For LLMs, these input–output pairs are most often Q&A examples used to evaluate assistants (Chang et al., 2023; Guo et al., 2023) or autonomously acting agents (Wang et al., 2024; Yehudai et al., 2025). The use of natural language gives a perspective of breadth, but fails to measure realistic human-LLM interaction (Guo et al., 2023) and hence real-world impact (Burden et al., 2025). This agrees with our findings that evaluations for data science mostly fail to capture intermediate levels of LLM-human collaboration and concentrate on evaluating substitution rather than higher levels of transformation (Sec. 2.1). Chang et al. (2023) highlighted human-in-the-loop testing and evaluations in an open environment as future directions, and Wang et al. (2024) identified a shift towards end-to-end tasks requiring human evaluators and versatile metrics, yet most evaluations today only consider a subset of tasks in the data science pipeline, with a few exceptions (Sec. 4.3).
A few studies evaluate truly long-horizon scenarios or quantify the human effort they still require. Wang et al. (2023) and Park et al. (2023) showed that agents can sustain hours-to-days of open-ended play or social simulation, but both exposed failure modes that need periodic human nudges. Quantitatively, Liu et al. (2023a) found that commercial models needed a median of 2.4 human corrections per task on a general agent benchmark, whereas open-source models needed 5–8. Recently, Kwa et al. (2025) showed that autonomous agents are progressively conquering tasks that take humans longer to complete when considering a fixed success rate (e.g., 50%), but performance still progressively degrades on tasks requiring more than 10 seconds.
@@ 2.4 Data science automation
Automating data science was a topic of research even before LLMs became commonplace. Bie et al. (2021) argued that the technical and domain knowledge required to solve data science tasks motivated efforts toward automation. The authors categorised data science tasks into four main quadrants, defined by two axes—degree of open-endedness and dependence on domain context—highlighting that model-building activities are more easily automated (e.g., through AutoML approaches, Hutter et al. 2019; Gijsbers et al. 2024) due to their lower open-endedness and context dependence. They also identified three forms of automation: mechanisation, composition, and assistance. Assistance corresponds to our interpretation of the term while mechanisation and composition can be grouped under our umbrella of automation (Sec. 2.2), but differ in focusing respectively on small parts of the process or on the overall pipeline; in our work, we do not make this distinction and instead rely on the activities of Martínez-Plumed et al. (2019) to identify how many elements of the pipeline each evaluation work covers. After Bie et al. (2021) published their survey, numerous works built LLM agents to automate data science. Their evolution, capabilities, and applications across the data science pipeline are reviewed by Sun et al. (2024); the authors, however, did not address LLM evaluation for data science, which is the focus of our work. More recently, (Hu et al., 2025) proposed a taxonomy for the data ecosystem: Data Management, which includes data collection, data storage, and data preprocessing; Data Analysis, which includes model evaluation, data interpretation, and decision making; and Data Visualisation. This taxonomy partly overlaps with that in Martínez-Plumed et al. (2019) which we use in our work (Sec. 2.5), but misses some of the most exploratory aspects.
@@ 2.5 The activities of data science
Table 1: Data-science activities and brief definition (complete definitions in Appendix A).
@@
Result Exploration (RE) Connect data-science results back to business goals
@@
Data Simulation (Sim) Simulate systems to generate data and explore causal “what-if” scenarios
@@
Many taxonomies of data science activities exist (see Martínez-Plumed et al., 2019, Sec 2). One of the most popular is CRISP-DM (Cross Industry Standard Process for Data Mining, Chapman, 2000), which considers projects as goal-oriented, with a pre-defined objective that can be approached by “mining” data through an approximately sequential process, from problem framing to solution delivery. However, Martínez-Plumed et al. (2019) argues that this goal-oriented, pre-collected-data perspective ignores many tasks of modern data science, where exploration is essential and the data takes centre stage rather than serving as a fixed backdrop. Consequently, Martínez-Plumed et al. (2019) proposes a sequence of exploratory activities that underscore the less prescriptive nature of data science, re-framing it as an investigative endeavour; and data management activities that treat data as dynamic rather than static. We provide a list of the activities introduced in Martínez-Plumed et al. (2019) and a concise definition in Table 1. Note that not all modern data-science projects include every activity, nor is the order of activities fixed as in the CRISP-DM framework. Instead, each project follows its own “trajectory” in the space of data-science tasks (Martínez-Plumed et al., 2019).
@@ 3 Evaluating LLM assistants in data science
In this section, we focus on evaluations of LLMs as assistants, namely prompting them in a fixed, pre-determined manner without letting them independently determine the sequence of steps. Table 2 shows the surveyed papers and the activities (Sec. 2.5) they cover; a double tick marks an activity that is explicitly assessed, whereas a single tick marks one that is vital for completing the tasks but not directly assessed.
@@
Goal-oriented Exploratory Data Management Papers BU DU DP M E Dep GE DSE DVE RE NE PE Acq Sim Arch Rel ARCADE (Yin et al., 2022) - ✓✓ - - - - - - - - - - - - CERT (Zan et al., 2022) - - ✓✓ - - - - - - - - - - - CoSQL (Yu et al., 2019a) - - - - ✓✓ ✓✓ - - - - - - - ✓✓ - DS-1000 (Lai et al., 2023) - ✓✓ ✓✓ - - - - - - - - - - - - DS-Bench (Ouyang et al., 2025) - ✓✓ ✓✓ - - - - - - - - - - - - DSP (Chandel et al., 2022) - ✓✓ - - - - - - - - - - - - - FeatEng (Pietruszka et al., 2024) - ✓✓ ✓✓ - - - - - - - - - - GPT4-DA (Cheng et al., 2023) ✓✓ ✓✓ - ✓✓ - - ✓✓ ✓✓ ✓✓ - - - - - HardML (Pricope, 2025) - - - - - - - - - - - LIDA (Dibia, 2023) - ✓✓ ✓✓ - - ✓✓ ✓✓ - ✓✓ - - - ✓✓ SParC (Yu et al., 2019b) - - - - - - - - - - - - - ✓✓ - Spider (Yu et al., 2018) - - - - - - - - - - - - - ✓✓ - Spider 2.0-Lite (Lei et al., 2024) - - - - - - - - - - - - - ✓✓ - Spider 2.0-Snow (Lei et al., 2024) - - - - - - - - - - - - - ✓✓ - StatLLM (Song et al., 2025) - ✓✓ ✓✓ - - - - - - - - - - -
@@
First, many works evaluate LLMs used to generate code for specific steps of data science, such as preprocessing data given a template, fixing bugs, or producing visualisations given instructions or prerequisites. In particular, ARCADE (Yin et al., 2022) and CERT (Zan et al., 2022) focus on Data Preparation and related activities with specific Python libraries. ARCADE is a benchmark consisting of 1,082 coding problems involving data wrangling and Exploratory Data Analysis (EDA), defined as Jupyter notebooks, that require Python’s Pandas library; for example, a problem could involve extracting min and max values from a dataframe and answer questions such as “In which year was the most played game added?”. CERT instead introduces two benchmarks (PandasEval and NumpyEval), each consisting of 101 tasks manually reworked for coherence and consistency from StackOverflow [^2] https://stackoverflow.com/questions problems tagged as relevant to Pandas and NumPy respectively; for problems whose solution is a function, 20 test cases are included, while the correctness of the predicted variable is checked for the other problems. Relatedly, DSP (Chandel et al., 2022) contains problems instantiated in 306 pedagogical Jupyter notebooks with 92 associated datasets, covering data manipulation, cleaning, and wrangling (parts of Data Preparation). Similarly to CERT, the correctness of the task is automatically graded with test cases. An example problem would be “Show the correlation between population density in 2023 and 2050, rounded to 2 decimals”. Similarly, DS-1000 (Lai et al., 2023) consists of 1,000 coding problems extracted from StackOverflow, spanning Python libraries such as NumPy, Pandas, SciKit-Learn, TensorFlow, matplotlib, SciPy, and PyTorch. The problems are manually perturbed to circumvent the issue of memorisation in LLMs and cover Data Preparation and Modelling. The problems are scored through multi-criteria execution-based evaluation metrics that rely on test cases and surface-form constraints to check for the presence of specific APIs. See Fig. 1 for an example problem and evaluation set-up. Ouyang et al. (2025) extend DS-1000 to create DS-Bench by adding Seaborn, Keras and LightGBM. To build the benchmark, they first define a broad task scope and collect Python seed code from GitHub using DS-1000’s reference code and corresponding StackOverflow answers. An automated LLM pipeline then transforms each snippet to avoid memorisation. Candidates are filtered by properties (compilability, stars, API calls) and functionality (must pass at least one LLM-generated test). For each surviving candidate, an LLM generates 200 test cases and a problem description—complete with an introduction, function signature, input/output formats and examples. After manual review, 1,000 problems are selected. Performance is measured with pass@$k$: a problem is solved if any $k$ generated samples pass the unit tests. Instead, FeatEng (Pietruszka et al., 2024), evaluates LLMs’ ability to produce a Python function for engineering data features (thus addressing Data Understanding and Preparation) suitable for downstream modelling tasks. The authors select datasets based on their popularity on Kaggle and ensuring broad domain coverage, reaching a total of 101 tasks. Notably, and in contrast to the older works described above, where questions admitted fixed ground truths, performance is measured in terms of the reduction in error of a model trained on the extracted features compared to a baseline model trained on the original, untransformed data. Therefore, this allows AI models to go beyond simply substituting humans and reach higher levels of task transformation (Sec. 2.1). StatLLM (Song et al., 2025) instead focuses on statistical Modelling, assessing LLMs’ ability to generate code to solve a dataset of 207 statistical analysis tasks assembled from various public online resources, including descriptive statistics, hypothesis testing, regression and ANOVA, generalised linear models, survival analysis, model selection, and non-parametric statistics; tasks might require the LLM to run a specific model on a variable in a given dataset, or to plot a variable. Uniquely, the LLM has to generate SAS code; evaluation is carried out using Natural Language Processing (NLP) metrics to compare the generated code against a human gold standard, thus being grounded in substitution (Sec. 2.1).

### v2

@@ Abstract
Data science aims to extract insights from data to support decision-making processes. Recently, Large Language Models (LLMs) have been increasingly used as assistants for data science, by suggesting ideas, techniques and small code snippets, or for the interpretation of results and reporting. Proper automation of some data-science activities is now promised by the rise of LLM agents, i.e., AI systems powered by an LLM equipped with additional affordances—such as code execution and knowledge bases—that can perform self-directed actions and interact with digital environments. In this paper, we survey the evaluation of LLM assistants and agents for data science. We find (1) a dominant focus on a small subset of goal-oriented activities, largely ignoring data management and exploratory activities; (2) a concentration on pure assistance or fully autonomous agents, without considering intermediate levels of human-AI collaboration; and (3) an emphasis on human substitution, therefore neglecting the possibility of higher levels of automation thanks to task transformation.
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020) and their multimodal extensions first caught public prominence by powering capable chatbots that are now widely used as assistants to humans in several tasks, such as summarising documents (Liu et al., 2023c), performing translations (Zhu et al., 2023), and creating code snippets (Guo et al., 2023). These LLM assistants take instructions from a human in the form of a prompt and return an answer, with the human retaining control over planning and decision-making by determining the sequence of actions to follow and deciding how much to rely on the assistant’s output; while they can use tools, their usage is generally prescribed at certain steps. Now, attention is increasingly dedicated to “LLM agents(Wang et al., 2024) that can autonomously and iteratively decide a sequence of actions to take and repeatedly interact with an external (digital) environment, being equipped with affordances and tools such as code execution (Huang et al., 2024b), internet access (Zhou et al., 2024), knowledge bases (Chen et al., 2024), and operating system control (Liu et al., 2023a), which the agents can decide to use of autonomously.
@@
In this paper we focus on how assistants and agents for data science applications are being evaluated. Our aim is to provide an overview of current evaluation set ups, and highlight the gaps in scopes of evaluation, rather than reviewing the performance of the models themselves. Data science is the process of handling and analysing data to extract insights that support decision-making in science, business, or other contexts. Data science may deal with different modalities of data (tabular, images, audio, etc.) but it always involves writing and interpreting information represented as text, whether in the form of code or natural language, and processing images, such as to present information or models. This dominance of textual modalities, combined with the vast amount of relevant online material that LLMs leverage during both training and inference, makes data science well-suited for automation by LLMs. Indeed, since the early days of LLMs (Chandel et al., 2022), LLM assistants and agents have been used in data science applications. In this work, we survey evaluation tools measuring performance on tasks across the data science pipeline, offering, to our knowledge, the first comprehensive review of LLM evaluation for core data science tasks [^1] While many studies assess LLMs on general coding tasks (Jimenez et al., 2024) and planning (Valmeekam et al., 2023), our survey focuses specifically on evaluations within the data science domain. This includes works that target data science-specific coding as well as other data science activities that are often overlooked in existing evaluations..
@@
Data science is highly multidisciplinary and involves a breadth of activities, combining fields such as statistics, machine learning, and data engineering with tasks such as understanding business needs, writing reports, and preliminary research. Therefore, to compare the data science automation evaluation tools in our survey, we adopt the data science task taxonomy of Martínez-Plumed et al. (2019), which expands the traditional CRISP-DM framework to include activities related to data management and exploratory analysis (see Fig. 1, reproducing Fig. 3 from Martínez-Plumed et al. 2019), and classify each evaluation tool by the activities it requires subjects to perform and explicitly evaluates. Moreover, in our survey we specifically consider the level of autonomy each evaluation targets—whether the subjects are agents, assistants, or intermediate forms, such as LLM agents operating under close human supervision that correct their actions as needed. Further, we also analyse the way in which the tasks are framed and evaluated, to understand if they measure the ability of AI systems to simply substitute humans or if instead they consider that the AI system can transform the task in deeper ways, such as by bringing functional improvements (for instance, LLMs may not need to produce visualisations to perform data exploration; see Sec. 2.1).
It is worth stressing how, in our work, we do not overview developments in state-of-the-art LLM assistants or agents for data science (referring interested readers to other works, Sun et al., 2024) nor analyse the current performance of such systems across the various data science activities and autonomy levels. While such efforts would be valuable and may reveal strong correlations between LLM systems’ performance on specific tasks, our work addresses a more foundational need: to establish taxonomies across which data science evaluations can be structured and to identify gaps in the evaluation landscape that must be filled before a comprehensive evaluation of LLM abilities can be carried out. Therefore, by focusing on activities and autonomy, we make the following findings:
@@
  • Most evaluation works individually target a (small) subset of data science activities (Tables 2 and 3); a few works cover multiple activities (Secs. 4.2 and 4.3), but none cover all of them. Taken together, the surveyed works cover the landscape of data‑science activities in a biased fashion, chiefly over‑representing “goal‑oriented” activities, such as data preprocessing, producing plots in specified formats, or building predictive models for predetermined targets. This prevents the identification of correlations between AI systems’ performance on different activities. Only a few studies (Cheng et al., 2023; Sahu et al., 2025; Majumder et al., 2024) give prominence to open-ended, exploratory aspects of data science (such as interpreting client needs within a business context, creatively exploring datasets and proposing potential uses) or data management(Yu et al., 2018; 2019b; 2019a; Lei et al., 2024).
@@
  • Most studies focus either on assistants following human-defined actions or fully autonomous agents, overlooking more realistic scenarios of intermediate LLM–human collaboration (also referred to as “centaur evaluations”, Haupt & Brynjolfsson, 2025). Exceptions include Li et al. (2025c; b; a) (agents with simulated human users) and Yu et al. (2019a) (assistants aiding users in clarifying data tasks). Due to this near-binary focus, we organise the surveyed works according to the focus on assistants (Sec. 3) and agents (Sec. 4), highlighting when they assess intermediate autonomy.
@@
  • More than half of the evaluations we survey implicitly assume that AI will substitute humans without functionally changing the tasks, either in assuming the steps by which a task is solved are the same a human would follow (Yu et al., 2019a; Zhang et al., 2024b; Chen et al., 2024) or by scoring the task referring to human-produced output, despite there not being a single ground truth (Song et al., 2025; Huang et al., 2024b; Hu et al., 2024; Jing et al., 2024; Chen et al., 2024). Examples of works also rewards agents that functionally transform the task to solve it by scoring the final output are Pietruszka et al. (2024); Cheng et al. (2023); Li et al. (2025b); Chan et al. (2024); Lei et al. (2024); Majumder et al. (2024); Sahu et al. (2025).
@@ 2.3 LLM evaluation
The area of AI evaluation (Burden et al., 2025) mostly relies on tasks encapsulated in input-output benchmarks with a reference output for each example. For LLMs, these input–output pairs are most often Q&A examples used to evaluate assistants (Chang et al., 2023; Guo et al., 2023) or autonomously acting agents (Wang et al., 2024; Yehudai et al., 2025). While the use of natural language gives a perspective of breadth, recent works highlighted how current evaluation practice fails to measure realistic human-LLM interaction (Haupt & Brynjolfsson, 2025) and real-world impact (Burden et al., 2025; Reiter, 2025). This agrees with our findings that evaluations for data science mostly fail to capture intermediate levels of LLM-human collaboration and concentrate on evaluating substitution rather than higher levels of transformation (Sec. 2.1). Indeed, Chang et al. (2023) and Haupt & Brynjolfsson (2025) highlighted human-in-the-loop testing and evaluations in an open environment as important directions, and Wang et al. (2024) identified a shift towards end-to-end tasks requiring human evaluators and versatile metrics, yet most evaluations today only consider a subset of tasks in the data science pipeline, with a few exceptions (Sec. 4.3).
A few studies evaluate truly long‑horizon scenarios or quantify the human effort they still require. Wang et al. (2023) and Park et al. (2023) showed that agents can sustain hours‑to‑days of open‑ended play or social simulation, but both exposed failure modes that need periodic human nudges. Quantitatively, Liu et al. (2023a) found that commercial models needed a median of 2.4 human corrections per task on a general agent benchmark, whereas open‑source models needed 5–8. Recently, Kwa et al. (2025) showed that autonomous agents are progressively conquering tasks that take humans longer to complete when considering a fixed success rate (e.g., 50%), but performance still progressively degrades on tasks requiring more than 10 seconds. Recently, Kwa et al. (2025) stratified tasks according to the mean execution time humans require to complete them, then measured AI model performance on each subset of tasks within specific time ranges (e.g., tasks taking around 10 minutes for humans). Plotting human-estimated task completion time against AI success rate, they demonstrated that AI success rates decrease as human completion time increases. However, when examining the human time range at which different AI models achieve a 50% success rate, more recent models consistently reach this threshold on tasks requiring longer human completion times, indicating progressive improvement in tackling more complex, time-intensive tasks.
@@ 2.4 Data science automation
Automating data science was a topic of research even before LLMs became commonplace. Bie et al. (2021) argued that the technical and domain knowledge required to solve data science tasks motivated efforts toward automation. The authors categorised data science tasks into four main quadrants, defined by two axes—degree of open-endedness and dependence on domain context—highlighting that model-building activities are more easily automated (e.g., through AutoML approaches, Hutter et al. 2019; Gijsbers et al. 2024) due to their lower open-endedness and context dependence. They also identified three forms of automation: mechanisation, composition, and assistance. Assistance corresponds to our interpretation of the term while mechanisation and composition can be grouped under our umbrella of automation (Sec. 2.2), but differ in focusing respectively on small parts of the process or on the overall pipeline; in our work, we do not make this distinction and instead rely on the activities of Martínez-Plumed et al. (2019) to identify how many elements of the pipeline each evaluation work covers. After Bie et al. (2021) published their survey, numerous works built LLM agents to automate data science. Their evolution, capabilities, and applications across the data science pipeline are reviewed by Sun et al. (2024); the authors, however, did not address LLM evaluation for data science, which is the focus of our work. More recently, (Hu et al., 2025) proposed a taxonomy for the data ecosystem: Data Management, which includes data collection, data storage, and data preprocessing; Data Analysis, which includes model evaluation, data interpretation, and decision making; and Data Visualisation. This taxonomy partly overlaps with that in Martínez-Plumed et al. (2019) which we use in our work (Sec. 2.5), but misses some of the most exploratory aspects. Finally, Chintakunta et al. (2025) conduct a systematic mapping study examining the application of LLMs in data science. Considering a 5-stage decomposition of data science, they find that most papers apply LLMs in Data Exploration and Analysis, followed by Model Building and Evaluation, and Data Collection and Preparation, with Deployment and Problem Definition unexplored – which mostly matches our findings, obtained with a finer task decomposition. Moreover, from the surveyed papers, a consensus on research gaps emerges, including the need to expand to complex, real-world evaluation tasks and to enhance human-AI interaction by creating more user-friendly interfaces; these echo our findings on the lack of human-AI interaction evaluations and the artificial nature of many evaluation tools, for instance by the adoption of a “substitution” perspective. In addition, our survey complements Chintakunta et al. (2025) by providing a deeper analysis of evaluation.
@@ 2.5 The activities of data science
Table 1: Data‑science activities and brief definition (complete definitions in Appendix A).
@@
Result Exploration (RE) Connect data‑science results back to business goals
@@
Data Simulation (Sim) Simulate systems to generate data and explore causal “what‑if” scenarios
@@
Many taxonomies of data science activities exist (see Martínez-Plumed et al., 2019, Sec 2). One of the most popular is CRISP-DM (Cross Industry Standard Process for Data Mining, Chapman, 2000), which considers projects as goal-oriented, with a pre-defined objective that can be approached by “mining” data through an approximately sequential process, from problem framing to solution delivery. However, Martínez-Plumed et al. (2019) argues that this goal‑oriented, pre‑collected‑data perspective ignores many tasks of modern data science, where exploration is essential and the data takes centre stage rather than serving as a fixed backdrop. Consequently, Martínez-Plumed et al. (2019) expands this taxonomy by proposing a sequence of exploratory activities that underscore the less prescriptive nature of data science, re-framing it as an investigative endeavour; and data management activities that do not assume data is already given and require to fetch more data from different sources. We provide a list of the activities introduced in Martínez-Plumed et al. (2019) and a concise definition in Table 1. Note that not all modern data-science projects include every activity, nor is the order of activities fixed as in the CRISP-DM framework. Instead, each project follows its own “trajectory” in the space of data-science tasks (Martínez-Plumed et al., 2019). Moreover, the distinction between activities may be blurred.
@@ 3 Evaluating LLM assistants in data science
In this section, we focus on evaluations of LLMs as assistants, namely prompting them in a fixed, pre-determined manner without letting them independently determine the sequence of steps. Table 2 shows the surveyed papers and the activities (Sec. 2.5) they cover; a double tick marks an activity that is explicitly assessed, whereas a single tick marks an activity that is required for completing the tasks but not directly assessed.
@@
Goal-oriented Exploratory Data Management Papers BU DU DP M E Dep GE DSE DVE RE NE PE Acq Sim Arch Rel ARCADE (Yin et al., 2022) - ✓✓ - - - - - - - - - - - - AssistedDS (Luo et al., 2025) - - - - - ✓✓ - - - - - - - CERT (Zan et al., 2022) - - ✓✓ - - - - - - - - - - - CoSQL (Yu et al., 2019a) - - - - ✓✓ ✓✓ - - - - - - - ✓✓ - DS-1000 (Lai et al., 2023) - ✓✓ ✓✓ - - - - - - - - - - - - DS-Bench (Ouyang et al., 2025) - ✓✓ ✓✓ - - - - - - - - - - - - DSP (Chandel et al., 2022) - ✓✓ - - - - - - - - - - - - - FeatEng (Pietruszka et al., 2024) - ✓✓ ✓✓ - - - - - - - - - - GPT4-DA (Cheng et al., 2023) ✓✓ ✓✓ - ✓✓ - - ✓✓ ✓✓ ✓✓ - - - - - HardML (Pricope, 2025) - - - - - - - - - - - LIDA (Dibia, 2023) - ✓✓ ✓✓ - - ✓✓ ✓✓ - ✓✓ - - - ✓✓ SParC (Yu et al., 2019b) - - - - - - - - - - - - - ✓✓ - Spider (Yu et al., 2018) - - - - - - - - - - - - - ✓✓ - Spider 2.0-Lite (Lei et al., 2024) - - - - - - - - - - - - - ✓✓ - Spider 2.0-Snow (Lei et al., 2024) - - - - - - - - - - - - - ✓✓ - StatLLM (Song et al., 2025) - ✓✓ ✓✓ - - - - - - - - - - -
@@
First, many works evaluate LLMs used to generate code for specific steps of data science, such as preprocessing data given a template, fixing bugs, or producing visualisations given instructions or prerequisites. In particular, ARCADE (Yin et al., 2022) and CERT (Zan et al., 2022) focus on Data Preparation and related activities with specific Python libraries. ARCADE is a benchmark consisting of 1,082 coding problems involving data wrangling and Exploratory Data Analysis (EDA), defined as Jupyter notebooks, that require Python’s Pandas library; for example, a problem could involve extracting min and max values from a dataframe and answer questions such as “In which year was the most played game added?”. CERT instead introduces two benchmarks (PandasEval and NumpyEval), each consisting of 101 tasks manually reworked for coherence and consistency from StackOverflow [^2] https://stackoverflow.com/questions problems tagged as relevant to Pandas and NumPy respectively; for problems whose solution is a function, 20 test cases are included, while the correctness of the predicted variable is checked for the other problems. Relatedly, DSP (Chandel et al., 2022) contains problems instantiated in 306 pedagogical Jupyter notebooks with 92 associated datasets, covering data manipulation, cleaning, and wrangling (parts of Data Preparation). Similarly to CERT, the correctness of the task is automatically graded with test cases. An example problem would be “Show the correlation between population density in 2023 and 2050, rounded to 2 decimals”. Similarly, DS-1000 (Lai et al., 2023) consists of 1,000 coding problems extracted from StackOverflow, spanning Python libraries such as NumPy, Pandas, SciKit-Learn, TensorFlow, matplotlib, SciPy, and PyTorch. The problems are manually perturbed to circumvent the issue of memorisation in LLMs and cover Data Preparation and Modelling. The problems are scored through multi-criteria execution-based evaluation metrics that rely on test cases and constraints to check whether the output relies on specific packages and functions. See Fig. 2 for an example problem and evaluation set-up. Ouyang et al. (2025) extend DS-1000 to create DS-Bench by adding Seaborn, Keras and LightGBM. To build the benchmark, they first define a broad task scope and collect Python seed code from GitHub using DS-1000’s reference code and corresponding StackOverflow answers. An automated LLM pipeline then transforms each snippet to avoid memorisation. Candidates are filtered by properties (compilability, stars, API calls) and functionality (must pass at least one LLM-generated test). For each surviving candidate, an LLM generates 200 test cases and a problem description—complete with an introduction, function signature, input/output formats and examples. After manual review, 1,000 problems are selected. Performance is measured with pass@$k$: a problem is solved if any $k$ generated samples pass the unit tests. Instead, FeatEng (Pietruszka et al., 2024), evaluates LLMs’ ability to produce a Python function for engineering data features (thus addressing Data Understanding and Preparation) suitable for downstream modelling tasks. The authors select datasets based on their popularity on Kaggle and ensuring broad domain coverage, reaching a total of 101 tasks. Notably, and in contrast to the older works described above, where questions admitted fixed ground truths, performance is measured in terms of the reduction in error of a model trained on the extracted features compared to a baseline model trained on the original, untransformed data. Therefore, this allows AI models to go beyond simply substituting humans and reach higher levels of task transformation (Sec. 2.1). StatLLM (Song et al., 2025) instead focuses on statistical Modelling, assessing LLMs’ ability to generate code to solve a dataset of 207 statistical analysis tasks assembled from various public online resources, including descriptive statistics, hypothesis testing, regression and ANOVA, generalised linear models, survival analysis, model selection, and non-parametric statistics; tasks might require the LLM to run a specific model on a variable in a given dataset, or to plot a variable. Uniquely, the LLM has to generate SAS code; evaluation is carried out using Natural Language Processing (NLP) metrics to compare the generated code against a human gold standard, thus being grounded in substitution (Sec. 2.1).